The area of a triangular sign is 33 square meters. the base of the triangle is 1meter less than double the altitude. find the altitude and the base of the sign
Answer by nerdybill(7384)

The area of a triangular sign is 33 square meters. the base of the triangle is 1meter less than double the altitude. find the altitude and the base of the sign
.
Let a = altitude
then
2a-1 = base
.
definition of area of any triangle:
(1/2)*base*height
.
(1/2)(2a-1)(a) = 33
(2a-1)(a) = 66
2a^2-a = 66
2a^2-a-66 = 0
(2a-11)(a+6) = 0
a = {11/2, -6}
we can throw out the negative solution leaving
a = 11/2 = 5.5 meters (altitude)
.
Base:
2a-1 = 2(5.5)-1 = 11-1 = 10 meters